OpenStudy (anonymous):

If given the slope m=10/3 and the points (1,-4) and (4,k) how do I solve for k

jimthompson5910 (jim_thompson5910):

do you know the slope formula?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

or \[\frac{ k-(-4) }{ 4-1 }=\frac{ 10 }{ 3 }\]